# Towards a Dynamic Future with Adaptable Computing and Network Convergence (ACNC)

### v1

@@ Abstract
In the context of advancing 6G, a substantial paradigm shift is anticipated, highlighting comprehensive everything-to-everything interactions characterized by numerous connections and stringent adherence to Quality of Service/Experience (QoS/E) prerequisites. The imminent challenge stems from resource scarcity, prompting a deliberate transition to Computing-Network Convergence (CNC) as an auspicious approach for joint resource orchestration. While CNC-based mechanisms have garnered attention, their effectiveness in realizing future services, particularly in use cases like the Metaverse, may encounter limitations due to the continually changing nature of users, services, and resources. Hence, this paper presents the concept of Adaptable CNC (ACNC) as an autonomous Machine Learning (ML)-aided mechanism crafted for the joint orchestration of computing and network resources, catering to dynamic and voluminous user requests with stringent requirements. ACNC encompasses two primary functionalities: state recognition and context detection. Given the intricate nature of the user-service-computing-network space, the paper employs dimension reduction to generate live, holistic, abstract system states in a hierarchical structure. To address the challenges posed by dynamic changes, Continual Learning (CL) is employed, classifying the system state into contexts controlled by dedicated ML agents, enabling them to operate efficiently. These two functionalities are intricately linked within a closed loop overseen by the End-to-End (E2E) orchestrator to allocate resources. The paper introduces the components of ACNC, proposes a Metaverse scenario to exemplify ACNC’s role in resource provisioning with Segment Routing v6 (SRv6), outlines ACNC’s workflow, details a numerical analysis for efficiency assessment, and concludes with discussions on relevant challenges and potential avenues for future research.
@@
6G, Deterministic Networking, Computing-Network Convergence (CNC), Joint Resource Allocation, the Metaverse, Dynamic, Autonomous Orchestration, Dimension Reduction, Context Detection, Continual Learning, Machine Learning, and Segment Routing v6 (SRv6).I Introduction
@@
In the foreseeable future for the 6th generation of communication systems (6G), interrelations are anticipated to enlarge beyond purely human-to-human connections to include everything-to-everything interactions. The Metaverse, a virtual realm of boundless possibilities, is poised for an era of rapid and explosive expansion. In this digital universe, a diverse multitude of both physical and digital objects will incessantly engage, traverse, and coexist within a seamlessly interconnected global network. This dynamic and immersive environment is expected to be further amplified by the persistent growth of connected objects, giving rise to a prolific exchange of data and generating substantial volumes of upstream traffic. In essence, it is expected that over billion entities [1] with device density of [2] and zettabytes of global data will be the norm of 6G [3]. This explosive surge in the demand should be acknowledged considering that up to 50% of this traffic will be required to be transmitted to computing nodes, processed, and returned to users [1].
@@
An additional aspect of the swift progress of technology pertains to the wide spectrum of promising and innovative services expected to materialize in the near future. One notable example is the concept of Metaverse as a Service (MaaS), empowering users to create, host, and manage personalized virtual environments across various domains like healthcare, finance, entertainment, and education. When combined with other capabilities, such as Spatial Computing, facilitating the convergence of physical and digital realms through Virtual, Augmented, and Extended Realities (VR/AR/XR) [4] to create a Digital Twin of a Person (DToP), these combined capabilities yield hybrid services replicating an individual’s distinctive attributes and enabling their coexistence in multiple locations, bridging both digital and physical domains simultaneously. It is crucial to emphasize that these services require strict adherence to Quality of Service/Experience (QoS/E) standards. Key Performance Indicators (KPIs) integral to their realization include achieving an End-to-End (E2E) delay under millisecond, a reliability rate surpassing %, and a peak data rate of Terabit per second (Tbps) [2].
@@
The primary challenge in implementing future services lies in addressing the substantial demands on computing and network resources, coupled with stringent QoS/E requirements, stemming from inherent resource scarcity. A promising strategy to effectively address this challenge involves shifting from the prevailing paradigm of ”user network computing” to a more integrated approach denoted as Computing-Network Convergence (CNC), symbolized by ”user network computing”. This paradigm shift has garnered significant attention in recent years. Król et al. [6] explored this concept through Compute-First Networking (CFN), applying Information-Centric Networking (ICN) principles to automatically perform traffic routing on network devices based on data object names. Two complementary approaches, Compute-Aware Networking (CAN) [7] and Computing Power Networking (CPN) [3], empower network devices to include computing nodes’ capabilities and load statuses in their decisions. Zhang et al. [8] proposed a comprehensive CNC architecture integrating resource allocation for both computing and network domains. Additionally, Albalawi et al. [9] investigated a concept similar to CNC, termed Integrated Network and Compute Allocation (INCA).
@@
In the pursuit of realizing future services, particularly in use cases like the Metaverse, the effectiveness of joint-orchestration strategies may face limitations. An extra complication is the dynamic nature of user requirements, seamless transitions between services in virtual realms, and frequent traversal between virtual and physical environments. In such a dynamic ecosystem, the live and scalable distribution of complex system states to network devices, coupled with resource allocation decision-making, becomes crucial. Existing approaches are unsuitable for this task due to significant convergence times, potentially compromising stringent QoS/E requirements amidst numerous dynamic users and services [3, 10, 11]. This paper addresses a gap in the existing literature by proposing a Machine Learning (ML)-based solution for the joint orchestration problem, considering the dynamic nature of user behavior, service characteristics, and infrastructure resources. Our proposed solution, ACNC, seamlessly integrates ML-based state recognition and context detection across resources, as well as domain and E2E orchestrators. Utilizing hierarchical state recognition, ML-based reduction techniques efficiently handle the complexity of extracting system states from the vast user-service-computing-network space for resource allocation. Through context detection, dynamicity is handled by categorizing reduced yet ubiquitous system states into predefined contexts, enabling resource allocation agents dedicated to each context to efficiently converge to optimal solutions over similar states while maintaining divergence from distinct states.
@@ II-A The Interwoven Cloud-Network Harmony
Traditional resource orchestration methodologies are insufficient for organizing such integrated infrastructures to execute such complex scenarios, as they typically operate within individual domains, neglecting interdependencies among diverse domains and resources. In contrast, CNC-based approaches adopt a holistic perspective, considering the web of interactions for optimal E2E orchestration and efficient resource utilization. In the holographic conference scenario, CNC-based orchestration can integrate network link status, strategically placing service instances on computing nodes to leverage shared links and minimize network utilization for high-bandwidth requests. Additionally, CNC-based methodologies can mitigate ripple effects from failures or limited functionality within one domain on others. For instance, in the event of congestion in allocated network links, CNC-based approaches can reposition service instances to open new, efficient networking opportunities, maintaining strict QoS/E requirements. This adaptive approach contrasts with conventional methods that attempt to navigate congested links and seek alternative but merely feasible paths to the same instance locations.
@@ II-C The Wisdom of Continuous Conetx-Awareness
The sole viable approach to contend with resource allocation in dynamic, complex systems necessitates proactive preparation for potential changes and prompt, contextually appropriate responses. The most promising solution involves applying ML techniques, empowering intelligent agents to continually acquire, adapt, and make real-time, state-driven decisions. However, in scenarios of exceptionally high dynamicity, the process of learning new states may impede or overwrite previously acquired knowledge, resulting in catastrophic forgetting. This phenomenon involves the loss of the ability to make accurate decisions about previous states, hindering performance when dealing with multiple states learned sequentially. CL aims to mitigate this issue by introducing the concept of context [12], where system states are constructed and categorized into finite classes, or contexts, each with its distinct training process. Agents of different contexts are then isolated and trained in parallel, facilitating optimal resource allocation when encountering their respective states without being perturbed by the others. For illustrative purposes, the existing service instances on computing nodes can be used to define the current context. Introducing new instances or migrating/removing existing ones facilitates the loading of a new context. For example, , and , where is context , is computing node , and is service intsance .
@@ III-A The Parties Involved
This paper revolves around the establishment of services involving three key parties. The first party, termed the infrastructure, encompasses the integrated computing and network resources of future 6G systems, along with their orchestration components. The second party is the service provider, responsible for registering services within the system. These services, illustrated by scenarios like the holographic meeting, may involve single or multiple instances that communicate according to a predefined sequence and data format outlined in the service’s function chaining map specified during registration. The ultimate participant in the system is the user, responsible for submitting requests to access registered services. Each request requires an E2E connection, establishing a link between the user and the relevant service instance(s). This connection manages the transmission of data from diverse sources (e.g., microphones, cameras, positioning sensors) to the instances and delivers processed data, such as the live-rendered holographic meeting scene, back to the user.
@@ III-B The Architecture of Adaptable Converged Orchestration
With the consideration of the involved parties, the core challenge involves the placement of registered service instances onto computing nodes while simultaneously establishing E2E paths that facilitate request connectivity while traversing the required service instances, adhering to stringent QoS/E criteria. The overarching objective is to address this challenge in order to accommodate the future’s dynamic and large-scale demands effectively. To meet this objective, we introduce a framework named ACNC, representing the integration of CL principles with the CNC concept, with two primary objectives: 1) optimizing infrastructure capacity utilization through joint orchestration to maximize the number of supported users, and 2) simplifying the complexity by categorizing it into discrete contexts, thereby enabling proactive readiness to tackle context-specific fluctuations. The proposed framework consists of various layers and components is depicted in Fig. 2 and defined in what follows.
@@ III-B1 Resources
In the ACNC framework, each network device’s primary function is the creation of short-term states. This involves meticulous processing of traffic across all ports, where metrics like traffic characteristics, request age of information, and QoS/E requirements are measured and collected for all requests. The short-term resource state at time slot is denoted as , with a spatial dimensionality represented as 111Let denotes the number of users, signifies the number of active services, represents the quantity of requests per user and service, indicates the number of ports, and stands for the quantity of metrics measured per request.. This state is then stored in the device’s memory bank. To comprehend causal relationships between resource allocation decisions and state transitions, short-term resource states for time slots are aggregated to form a long-term resource state at time slot (). Given the complexity and multidimensionality of this state (), we employ ML techniques (such as Uniform Manifold Approximation and Projection (UMAP), which is known for its ability to handle large datasets) to reduce its dimensions. The resulting reduced resource state is denoted as , with size , where represents the size of the set representing each service over each port that can accommodate from one indicator to all metrics. A smaller set of indicators yields more concise reduced states, simplifying state exchanges but diminishing precision in state representations. After reduction, the device stores the reduced resource state in the memory bank and then forwards it to the network orchestrator. Given the reduced resource state size, the required time and resources for this exchange are predictable. The same operations are replicated on computing nodes, but they construct states for each service instance implemented on their infrastructure.
@@
For each edge network device or Point of Arrival (PoA), there is an additional task - the selection of service instances for incoming requests and determining their respective network paths. This task is executed in a context-aware manner, where each PoA is equipped with an agent dedicated to a unique context. These contexts are structured to enable a singular agent to proficiently oversee the dynamics of the system states associated with that context, leading to a notable reduction in overall complexity to a more manageable scale. The agent for each context maintains two action sets: 1) the available service instances within that context, and 2) the available network paths from the PoA to these service instances. When the context changes or a significant alteration occurs in the system state, the PoA receives relevant information about the active context and current system state in graph form (as detailed in Section III-B3) from the network orchestrator. Subsequently, the PoA activates the agent associated with the active context, assigning service instances and network paths to requests according to their chaining map within the specified time slot. Operating within a Reinforcement Learning (RL) framework for continuous reward-based training [13], this agent is reinforced by Graph Neural Networks (GNNs) utilizing a node-link representation to gain a deeper understanding of the system’s intricacies.
@@ III-B2 Domain Orchestrators
The orchestration layer consists of two key components: the network and computing orchestrators. Both orchestrators aggregate reduced resource states from members in their respective domains, establishing a comprehensive short-term domain state for each time slot (). The network orchestrator constructs a graphical representation for each service, reflecting the physical network’s port-level structure. Simultaneously, the computing orchestrator formulates an overlay map for each service based on its service instance chain. Moreover, the orchestrators monitor resource availability by collecting metrics such as operational uptime, failure events, resource utilization rates, power supply statuses, and environmental conditions, adding them to , with the number of metrics denoted by . is then stored in the memory bank. Now, aggregating short-term domain states for time slots, both orchestrators employ dimension reduction ML models to create reduced domain states () and share them with the E2E orchestrator. During this reduction, the domain state may be downsized from to 222Let denotes the number of network devices in the network domain or the number of service instances in the computing domain, signifies the number of network links, and stands for the quantity of metrics measured per resource., where represents the size of the domain-level service indicators set. to indicators may be included in this set, affecting the size and precision of reduced states. The final task involves conveying instructions from the E2E orchestrator to underlying resources, adhering to specific standards and data models, facilitated through the configuration functionality.

### v2

@@ Abstract
In the context of advancing 6G, a substantial paradigm shift is anticipated, highlighting comprehensive everything-to-everything interactions characterized by numerous connections and stringent adherence to Quality of Service/Experience (QoS/E) prerequisites. The imminent challenge stems from resource scarcity, prompting a deliberate transition to Computing-Network Convergence (CNC) as an auspicious approach for joint resource orchestration. While CNC-based mechanisms have garnered attention, their effectiveness in realizing future services, particularly in use cases like the Metaverse, may encounter limitations due to the continually changing nature of users, services, and resources. Hence, this paper presents the concept of Adaptable CNC (ACNC) as an autonomous Machine Learning (ML)-aided framework crafted for the joint orchestration of computing and network resources, catering to dynamic and voluminous user requests with stringent requirements. ACNC encompasses two primary functionalities: state recognition and context detection. Given the intricate nature of the user-service-resource space, the paper employs dimension reduction to generate live, holistic, abstract system states in a hierarchical structure. To address the challenges posed by dynamic changes, Continual Learning (CL) is employed, classifying the system state into contexts controlled by dedicated ML agents, enabling them to operate efficiently. These two functionalities are intricately linked within a closed loop overseen by the End-to-End (E2E) orchestrator to allocate resources. The paper introduces the components of ACNC, proposes a Metaverse scenario to exemplify ACNC’s role in resource provisioning with Segment Routing v6 (SRv6), outlines ACNC’s workflow, details a numerical analysis for efficiency assessment, and concludes with discussions on relevant challenges and potential avenues for future research.
@@
6G, Deterministic Networking, Computing-Network Convergence (CNC), Joint Resource Allocation, Metaverse, Autonomous Orchestration, Dimension Reduction, Context Detection, Continual Learning, Machine Learning, and Segment Routing v6 (SRv6).I Introduction
@@
In the foreseeable future for the 6th generation of communication systems (6G), interrelations are anticipated to enlarge beyond purely human-to-human connections to include everything-to-everything interactions. The Metaverse, a virtual realm of boundless possibilities, is poised for an era of rapid and explosive expansion. In this digital universe, a diverse multitude of both physical and digital objects will incessantly engage, traverse, and coexist within a seamlessly interconnected global network. This dynamic and immersive environment is expected to be further amplified by the persistent growth of connected objects, giving rise to a prolific exchange of data and generating substantial volumes of upstream traffic. In essence, the presence of billions of entities operating at exceptionally high densities, along with the generation of zettabytes of global data, is expected to be the norm of 6G [1]. This explosive growth in demand must be acknowledged, particularly as a substantial portion of this traffic will need to be transmitted to computing nodes for processing before being returned to users.
@@
An additional aspect of the swift progress of technology pertains to the wide spectrum of promising and innovative services expected to materialize in the near future. One notable example is the concept of Metaverse as a Service (MaaS), empowering users to create, host, and manage personalized virtual environments across various domains like healthcare, finance, entertainment, and education. When combined with other capabilities, such as Spatial Computing, facilitating the convergence of physical and digital realms through Virtual, Augmented, and Extended Realities (VR/AR/XR) [2] to create a Digital Twin of a Person (DToP), these combined capabilities yield hybrid services replicating an individual’s distinctive attributes and enabling their coexistence in multiple locations, bridging both digital and physical domains simultaneously. It is essential to emphasize that these services necessitate strict adherence to Quality of Service/Experience (QoS/E) standards. The realization of these services hinges on several Key Performance Indicators (KPIs), which include achieving millisecond-level End-to-End (E2E) latency, an ultra-high reliability rate, and a peak data rate in the terabit range.
@@
The primary challenge in realizing the envisioned future lies in addressing the substantial demands on computing and network resources, coupled with stringent QoS/E requirements, stemming from inherent resource scarcity. A promising strategy to effectively address this challenge involves shifting from the prevailing paradigm of ”user network computing” to a more integrated approach denoted as Computing-Network Convergence (CNC), symbolized by ”user network computing”. This paradigm shift has garnered significant attention in recent years. Król et al. [4] explored this concept through Compute-First Networking (CFN), applying Information-Centric Networking (ICN) principles to automatically perform traffic routing on network devices based on data object names. Two complementary approaches, Compute-Aware Networking (CAN) [5] and Computing Power Networking (CPN) [1], empower network devices to include computing nodes’ capabilities and load statuses in their decisions. Zhang et al. [6] proposed a comprehensive CNC architecture integrating resource allocation for both computing and network domains. Additionally, Albalawi et al. [7] investigated a concept similar to CNC, termed Integrated Network and Compute Allocation (INCA).
@@
The effectiveness of the mentioned joint-orchestration strategies may face limitations, particularly in use cases like the Metaverse. An extra complication is the dynamic nature of user requirements, seamless transitions between services in virtual realms, and frequent traversal between virtual and physical environments. In such a dynamic ecosystem, the live and scalable distribution of complex system states to network devices, coupled with resource allocation decision-making, becomes crucial. Existing approaches are unsuitable for this task due to significant convergence times, potentially compromising stringent QoS/E requirements amidst numerous dynamic users and services [1, 8, 9]. This paper addresses a gap in the existing literature by proposing a Machine Learning (ML)-based solution for the joint orchestration problem, considering the dynamic nature of user behavior, service characteristics, and infrastructure resources. Our proposed solution, Adaptable CNC (ACNC), seamlessly integrates ML-based state recognition and context detection across resources, as well as domain and E2E orchestrators. Utilizing hierarchical state recognition, ML-based reduction techniques efficiently handle the complexity of extracting system states from the vast user-service-infrastructure space for resource allocation. Through context detection, dynamicity is handled by categorizing reduced yet ubiquitous system states into predefined contexts, enabling resource allocation agents dedicated to each context to efficiently converge to optimal solutions over similar states while maintaining divergence from distinct states.
@@ II-A The Interwoven Cloud-Network Harmony
Traditional resource orchestration methodologies are insufficient for organizing such integrated infrastructures to execute such complex scenarios, as they typically operate within individual domains, neglecting interdependencies among diverse domains and resources. In contrast, CNC-based approaches adopt a holistic perspective, considering the web of interactions for optimal E2E orchestration and efficient resource utilization. In the holographic conference scenario, CNC-based orchestration can integrate network link status, strategically placing service instances on computing nodes to leverage shared links and minimize network utilization for high-bandwidth requests. Additionally, they can mitigate ripple effects from failures or limited functionality within one domain on the other. For instance, in the event of congestion in allocated network links, CNC-based approaches can reposition service instances to open new, efficient networking opportunities, maintaining strict QoS/E requirements. This adaptive approach contrasts with conventional methods that attempt to navigate congested links and seek alternative but merely feasible paths to the same instance locations.
@@ II-C The Wisdom of Continuous Conetx-Awareness
Efficient resource allocation in dynamic, complex systems requires adaptive strategies that can anticipate changes and respond appropriately in real-time. Traditional ML approaches frequently face challenges in such environments due to their inability to effectively manage non-stationary data. The acquisition of new information can undermine previously learned knowledge, leading to decreased performance on earlier tasks, a phenomenon known as catastrophic forgetting. CL presents a promising solution by enabling intelligent agents to continuously acquire new knowledge while retaining previously learned information, ensuring accurate decision-making across diverse system states [10]. To facilitate this, contexts are introduced in CL, which partition various system states into distinct categories, each representing a specific configuration or operational condition. By organizing system states into finite contexts with dedicated training processes, agents can specialize in particular scenarios without interference, thereby mitigating catastrophic forgetting and enhancing adaptability and responsiveness. Furthermore, the modularity of this approach allows agents to be trained in parallel across different contexts, utilizing parallel computing resources to reduce training time. In this paper, we propose that the current context can be defined by the existing service instances on computing nodes. The introduction of new instances or the migration or removal of existing ones facilitates the transition to a new context. For example, and , where denotes context , represents computing node , and indicates service instance .
@@ III-A The Parties Involved
This paper involves three key parties. The first party, termed the infrastructure, encompasses the integrated computing and network resources of future 6G systems, along with their orchestration components. We assume the infrastructure relies on established protocols and data formats, ensures that data exchanged between different systems are secured and accurately interpreted. The second party is the service provider, responsible for registering services within the system. These services, illustrated by scenarios like the holographic meeting, may involve single or multiple instances that communicate according to a predefined sequence and data format outlined in the service’s function chaining map specified during registration. The ultimate participant in the system is the user, responsible for submitting requests to access registered services. Each request requires an E2E connection, establishing a link between the user and the relevant service instance(s). This connection manages the transmission of data from diverse sources (e.g., microphones, cameras, positioning sensors) to the instances and delivers processed data, such as the live-rendered holographic meeting scene, back to the user.
@@ III-B The Architecture of Adaptable Converged Orchestration
With the consideration of the involved parties, the core challenge involves the placement of registered service instances onto computing nodes while simultaneously establishing E2E paths that facilitate request connectivity and traversing the required service instances, adhering to stringent QoS/E criteria. The overarching objective is to address this challenge in order to accommodate the systems’s dynamic nature. To meet this, we introduce a framework named ACNC, representing the integration of CL principles with the CNC concept, with two primary principles: 1) optimizing infrastructure capacity utilization through joint orchestration to maximize the number of supported users, and 2) simplifying the complexity by defining discrete contexts, thereby enabling proactive readiness to tackle context-specific fluctuations. The proposed framework consists of various layers and components is depicted in Fig. 2 and defined in what follows.
@@ III-B1 Resources
In the ACNC framework, each network device’s primary function is the creation of short-term states. This involves meticulous processing of traffic across all ports, where metrics like traffic characteristics, request age of information, and QoS/E requirements are measured and collected for all requests. The short-term resource state at time slot is denoted as , with a spatial dimensionality represented as 111Let denotes the number of users, signifies the number of active services, represents the quantity of requests per user and service, indicates the number of ports, and stands for the quantity of metrics measured per request.. To comprehend causal relationships between resource allocation decisions and state transitions, short-term resource states for time slots are aggregated to form a long-term resource state at time slot (). Given the complexity and multidimensionality of this state (), we utilize ML techniques, including Uniform Manifold Approximation and Projection (UMAP), a method known for its ability to handle large datasets, to reduce dimensions. The resulting reduced resource state is denoted as , with size , where represents the size of the set representing each service over each port that can accommodate from one indicator to all metrics. A smaller set of indicators yields more concise reduced states, simplifying state exchanges but diminishing precision in state representations. After reduction, the device stores the reduced resource state in the memory bank and then forwards it to the network orchestrator. Given the reduced resource state size, the required time and resources for this exchange are predictable. The same operations are replicated on computing nodes, but they construct states for each service instance implemented on their infrastructure.
@@
For each edge network device or Point of Arrival (PoA), there is an additional task - the selection of service instances for incoming requests and determining their respective network paths. This task is executed in a context-aware manner, where each PoA is equipped with an agent dedicated to a unique context. These contexts are structured to enable a singular agent to proficiently oversee the dynamics of the system states associated with that context, leading to a notable reduction in overall complexity to a more manageable scale. The agent for each context maintains two action sets: 1) the available service instances within that context, and 2) the available network paths from the PoA to these service instances. When the context changes or a significant alteration occurs in the system state, the PoA receives relevant information about the active context and current system state in graph form (as detailed in Section III-B3) from the network orchestrator. Subsequently, the PoA activates the agent associated with the active context, assigning service instances and network paths to requests according to their chaining map within the specified time slot. We employ a Reinforcement Learning (RL) framework to implement this agent, wherein the system state and request information serve as inputs. The agent takes actions (instances and paths) to maximize a cumulative reward signal over time (see Section III-B3). To process the input effectively, the agent is implemented using Graph Neural Networks (GNNs) [11]. The GNN aggregates information from neighboring nodes within the system state graph, allowing the agent to incorporate both local context and global state conditions into its decision-making process.
@@ III-B2 Domain Orchestrators
The orchestration layer consists of two key components: the network and computing orchestrators. Both orchestrators aggregate reduced resource states from members in their respective domains, establishing a comprehensive short-term domain state for each time slot (). The network orchestrator constructs a graphical representation for each service, reflecting the physical network’s port-level structure. Simultaneously, the computing orchestrator formulates an overlay map for each service based on its service instance chain. Moreover, the orchestrators monitor resource availability by collecting metrics such as operational uptime, failure events, resource utilization rates, power supply statuses, and environmental conditions, adding them to , with the number of metrics denoted by . Now, aggregating short-term domain states for time slots, both orchestrators employ dimension reduction ML models to create reduced domain states () and share them with the E2E orchestrator after storing them. During this reduction, the domain state may be downsized from to 222Let denotes the number of network devices in the network domain or the number of service instances in the computing domain, signifies the number of network links, and stands for the quantity of metrics measured per resource., where represents the size of the domain-level service indicators set. to indicators may be included in this set, affecting the size and precision of reduced states. The final task involves conveying instructions from the E2E orchestrator to underlying resources, adhering to specific standards and data models, facilitated through the configuration functionality.